OpenStudy (anonymous):

Determine whether u and v are orthogonal, parallel, or neither. u=<1,2,-3> v=<(-4/3), (-8/3), 4> I dont understand how to do this. Can someone please help?

ganeshie8 (ganeshie8):

hint : two vectors are parallel if the components form same ratio

OpenStudy (anonymous):

What does that mean? I am sorry, I am not that great when it comes to vectors.

ganeshie8 (ganeshie8):

take the ratio of each component

ganeshie8 (ganeshie8):

ratio of x components : 1/(-4/3) = ? ratio of y components : 2/(-8/3) = ? ratio of z components : -3/4 = ?

ganeshie8 (ganeshie8):

if u get same number for all of them, then the vectors are parallel

OpenStudy (turingtest):

do you know how to take the dot product?

OpenStudy (anonymous):

The vectors are parallel. They all equaled the same number. Can you explain to me why they are parallel? Thanks

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

if the vectors are parallel, then we can write \(\large \vec{v} = t \vec{u}\) t is a scalar

ganeshie8 (ganeshie8):

right ?

ganeshie8 (ganeshie8):

say \(\large \vec{v} = <a_1, a_2, a_3>\), \(\large \vec{u} = <b_1, b_2, b_3>\) for them to be parallel : \(\large \vec{v} = t \vec{u} \) => \(\large <a_1, a_2, a_3> = t<b_1, b_2, b_3>\) \(\large <a_1, a_2, a_3> = <tb_1, tb_2, tb_3>\) compare the components both sides \(\large a_1 = tb_1\) \(\large a_2 = tb_2\) \(\large a_3 = tb_3\)

ganeshie8 (ganeshie8):

^^if u take the ratio of components u wud get the same number \(t\)

OpenStudy (anonymous):

So they are parallel then? Thats what I have been getting.

ganeshie8 (ganeshie8):

yup!

OpenStudy (anonymous):

Alright thanks! I really appreciate your help!

ganeshie8 (ganeshie8):

np :) see if u can reply to turingtest's question above... on dot product..